# Efficient Network Automatic Relevance Determination

### v1

@@ 1.2 Related work
$\left(\hat{W},\hat{\Omega}\right)=\underset{(W,\Omega)}{\operatorname{argmin}}\left\{l(W,\Omega)+\lambda_{1}\sum_{k\neq\ell}\left|\omega_{k\ell}\right|+\lambda_{2}\sum_{j=1}^{md}\left|w_{j}\right|\right\},$ (4)
@@
Graphical Lasso. The Graphical Lasso (GLasso) aims to estimate a sparse precision matrix $\Theta$ for a given dataset (Friedman et al., 2008). The general form of GLasso is defined as:
@@ 2.1 Framework of NARD
$\begin{split}W&\sim\mathcal{MN}(0,V_{m\times m},K^{-1}_{d\times d})\\ &=\frac{|K|^{m/2}}{(2\pi)^{md/2}|V|^{d/2}}\times\exp\left[-\frac{1}{2}\textbf{Tr}(V^{-1}WKW^{\top})\right].\end{split}$ (6)
@@ 2.3 Algorithm summary
[…]
@@ 3.2 Fast optimization of evidence
Thus, by rewriting $C$ as $C=C_{\backslash i}+\alpha_{i}^{-1}\varphi_{i}\varphi_{i}^{\top}$, and using determinant and inverse lemma of matrix, the logarithm of MLF can be explicitly decomposed into two parts, one part denoted by $\mathcal{L}(\alpha_{\backslash i})$, that does not depend on $\alpha_{i}$ and another that does, i.e.,
@@ 6.1 Synthetic datasets
Method $d$ $m$ $N$ TPR FPR Time (Total)
@@
Method $d$ $m$ $N$ TPR FPR Time (1 step)
@@ 6.2 Baselines
We further evaluate the scalability of NARD on larger datasets. As shown in 3, the experimental results align with the theoretical complexity, demonstrating that as the data size increases, the time per iteration for NARD grows in a manner consistent with its expected computational behavior. The Surrogate NARD approach consistently outperforms NARD, showing substantial improvements in computational efficiency, especially at larger scales.
@@ Appendix B Derivations in Detail
As a result, the first term in Eq. 9 can be decomposed into two parts: the first part corresponds to the kernel of the matrix normal distribution for $W$, and the second part corresponds to the kernel for $Y$.
@@
B.6 Details of Eq. 23
Eq. 21, Eq. 22 and Eq. 23 is the standard quantity in sparse learning. To some extent, $s_{i}$ is called the sparsity and $q_{i}$ is known as the quality of $\phi$, The sparsity measures the extent to which basis function overlaps with the other basis vectors in the model, and the quality represents a measure of the alignment of the basis vector with the error between the training set values and the vector of predictions that would result from the model with the vector excluded.
@@
In the proof of Eq. 21 below. We omit the subscript $\mathcal{A}$ presented in $C_{\mathcal{A}}$ for readability.
@@ C.2 Analysis of prior
When $\alpha_{i}\to\infty$ and fix $\alpha_{\backslash i}$, $|K|\to\infty$ then $A\to\infty$.
@@ D.2 Hyperprior for $\alpha$
Flat prior. If $\alpha$ follows the flat prior, i.e., $p(\alpha)=1$ then

### v2

@@ 1.2 Related work
$\small\left(\hat{W},\hat{\Omega}\right)=\underset{(W,\Omega)}{\operatorname{argmin}}\left\{l(W,\Omega)+\lambda_{1}\sum_{k\neq\ell}\left|\omega_{k\ell}\right|+\lambda_{2}\sum_{j=1}^{md}\left|w_{j}\right|\right\},$ (4)
@@
Graphical Lasso. The Graphical Lasso (GLasso) aims to estimate a sparse precision matrix $\Theta$ for a given dataset (Friedman et al., 2008; Banerjee et al., 2008; Yuan & Lin, 2007). The general form of GLasso is defined as:
@@ 2.1 Framework of NARD
$\small\begin{split}W&\sim\mathcal{MN}(0,V_{m\times m},K^{-1}_{d\times d})\\ &=\frac{|K|^{m/2}}{(2\pi)^{md/2}|V|^{d/2}}\times\exp\left[-\frac{1}{2}\textbf{Tr}(V^{-1}WKW^{\top})\right].\end{split}$ (6)
@@ 2.3 Algorithm summary
We can incorporate hyperpriors for the parameters $V$ and $K$ in our model. The hierarchical Bayesian approach allows for greater flexibility and robustness by capturing uncertainties in the prior distributions. We have $p(W;V)=\int p(W|K;V)p(K)d\alpha$. If each $\alpha_{i}$ follows a common Gamma distribution, then the prior of $W_{i}$ is a multidimensional Student distribution.
@@ 3.2 Fast optimization of evidence
[…]
@@ 6.1 Synthetic datasets
Method $d$ $m$ NN TPR FPR Time (Total)
@@
Method $d$ $m$ NN TPR FPR Time (1 step)
@@ 6.2 Baselines
We further evaluate the scalability of NARD on larger datasets. As shown in Table 3, the experimental results align with the theoretical complexity, demonstrating that as the data size increases, the time per iteration for NARD grows in a manner consistent with its expected computational behavior. The Surrogate NARD approach consistently outperforms NARD, showing substantial improvements in computational efficiency, especially at larger scales.
@@ Appendix B Derivations in Detail
As a result, the first term in Eq. (9) can be decomposed into two parts: the first part corresponds to the kernel of the matrix normal distribution for $W$, and the second part corresponds to the kernel for $Y$.
@@
B.6 Details of Eq. (23)
Eq. (21), Eq. (22) and Eq. (23) is the standard quantity in sparse learning. To some extent, $s_{i}$ is called the sparsity and $q_{i}$ is known as the quality of $\phi$, The sparsity measures the extent to which basis function overlaps with the other basis vectors in the model, and the quality represents a measure of the alignment of the basis vector with the error between the training set values and the vector of predictions that would result from the model with the vector excluded.
@@
In the proof of Eq. (21) below. We omit the subscript $\mathcal{A}$ presented in $C_{\mathcal{A}}$ for readability.
@@ C.2 Analysis of prior
[…]
@@ D.2 Hyperprior for $\alpha$
Flat prior. If $\alpha$ follows the flat prior, i.e., $p(\alpha)=1$, then